# README: Bridging Medical Jargon and Lay Understanding for Patient Education through Data-Centric NLP

### v1

@@ 4.2.4 Data Quality
The final stage involved preparing the training data, which comprised good definitions from README-v2 and README-v5. This data was split into three categories: human examination data (also be used as final test data since medical experts examine this split), out-of-distribution (OOD) evaluation data, and training data, where we make sure the medical jargon in human examination and OOD evaluation will not appear in the training split. The human examination split consisted of 1000 unique terms, each accompanied by general definitions to be rated based on two criteria:
  1. [topsep=0.5pt,itemsep=0.2ex,partopsep=0.2ex,parsep=.20ex]
@@
Although the lay definition is not incorrect, it is very wordy and complex and matches less to the lay definition. We consider this soft correlated and not hard correlated.
A Hard Correlation automatically implies a Soft Correlation. Our findings revealed that 88% of README-v2 and 100% of README-v5 met the Hard Correlation criteria. There is a 100% soft Correlation criteria for both README-v2 and README-v5.
4.3 Integration of Synthetic and Expert Data
We adopted four distinct sampling strategies to integrate the AI-synthetic README-v5 data with the expert-annotated README-v2 dataset:
  • [leftmargin=.2in,topsep=2pt]
  • RANDOM: This approach randomly selected N entries from the README-v5 dataset. This is the baseline for our subsequent three heuristic methods.
  • SEMANTIC: For semantic-based sampling, we utilized the all-MiniLM-L6-v2 model from the SentenceTransformers framework  [^3] https://github.com/UKPLab/sentence-transformers. Renowned for its text semantic analysis efficiency, this model enabled us to measure the semantic similarity between lay definitions in README-v2 and README-v5 datasets. We ranked synthetic data based on these scores, considering higher scores as indicative of greater semantic closeness to expert annotations.
  • MODEL: In model-based sampling, we used models initially trained on the README-v2 dataset to generate definitions for the README-v5 dataset. This tested the model’s capacity to transfer learnings from expert data to synthetic data. We employed the ROUGE_L F1 score to evaluate the alignment between model-generated and actual README-v5 lay definitions. This technique aids in mitigating training challenges associated with data heterogeneity and enriches the dataset with examples that enhance the model’s convergence towards the desired distribution (e.g., expert-annotated lay definitions).
5 Experiments
5.1 Automatic Evaluation Metrics
Models are evaluated with full-length F1-scores of ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005). We use QuickUMLS [^4] https://github.com/Georgetown-IR-Lab/QuickUMLS to extract medical concepts from both model-generated and ground truth summaries and then calculate F1-scores for these two lists of concepts, which is named UMLS-F1 Adams et al. (2023).
5.2 Experimental Setting
We use the following symbols:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
We use the following base models: GPT-2  [^5] https://huggingface.co/gpt2 Radford et al. (2019), DistilGPT2  [^6] https://huggingface.co/distilgpt2, BioGPT  [^7] https://huggingface.co/microsoft/biogpt Luo et al. (2022), and Llama2  [^8] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023) in our experiments. We trained the base models on the different README dataset variants with Supervised Fine-tuning for 100000 steps (batch size 8)  [^9] We did all the experiments with 1 NVIDIA Tesla RTX 8000 GPU - 40 GB memory, with Adam optimizer – betas=(0.9,0.999), epsilon=1e-08, learning rate=5e-04.. In all our evaluations, we used a beam size of 4, no-repeat-ngram-size=2, and minimum length and maximum length of sentences were set as (10, 100). We used five different random seeds to sample training data for all our experiments, and the scores reported in the tables are the average of these random seeds.
@@ 5.2 Experimental Setting
Our experiments were divided into five distinct parts. In Set-1, we aimed to evaluate the performance gap on the jargon2lay task between open-source base models and GPT models. We do one-shot prompting for Llama2, GPT-3.5-turbo, and GPT-4 with prompts in Table 1. Set-2 explored the varying data quality across different versions within our EAE pipeline, where we fine-tuned the GPT-2 model on the jargon+gen2lay task. Set-3 focused on evaluating the effects of various data selection strategies on data augmentation outcomes. To do this, we ranked AI-generated data (v5) using different methods in Section 4.3, selecting either the top N entries with the highest scores (e.g., ‘SEMANTIC’ in Table 3) or the bottom N entries with the lowest scores (e.g., ‘SEMANTIC_r’). These selections were then evenly combined with expert-annotated data (v2) at a one-to-one ratio. Following this, we fine-tuned the GPT-2 model using these diverse, mixed datasets to determine the impact of each selection method on model performance. In Set-4, we investigated the effects of incorporating different types of information (medical documentation context and UMLS-retrieved general definition) into the model inputs. In Set-5, we fine-tuned models of varying sizes (ranging from DistilGPT2-88M to Llama2-7B) and compared the outcomes with GPT-3.5-turbo. The next section (Section 6) further validates the findings obtained through this section’s automatic evaluation with a comprehensive human evaluation approach.
5.3 Results
We start with our experiments with some base models’ performance on the README jargon2lay dataset with In Set-1. In Section 4.2.4, we noted a high level of agreement between human evaluators and definitions generated by GPT-3.5-turbo. This aligns with the findings of recent work by Remy and Demeester (2023), which underlines GPT-3.5-turbo’s efficacy in crafting human-readable explanations for biomedical concepts. Consequently, GPT-3.5-turbo serves as a strong baseline for quality in README lay definition generation, indicative of a standard that meets human satisfaction. Despite this, our assessments of Llama2, depicted in Figure 4, show a notable gap in performance. This disparity underscores the urgent need to upgrade the performance of Llama2 and other open-source models (e.g., DistilGPT2, GPT2, BioGPT) to align with the high-quality output of GPT-3.5-turbo. Finally, We also observed that GPT-3.5-turbo and GPT4 performed equally well in this task. Next, we discuss how different modules of our data-centric pipeline improve the performance of the fine-tuned models in several subsections.
Figure 4: One-shot performances on jargon2lay.
5.3.1 Effectiveness of EAE pipeline
J+G2L ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1 Rank
v1 23.94 8.95 22.79 17.75 12.83 2
v2 26.99 10.76 25.57 21.19 16.88 4
v3 22.25 7.46 21.09 17.18 11.72 5
v5 25.71 9.75 24.41 20.23 16.01 3
v2+v5 29.82 13.14 28.47 24.42 20.09 1
Table 2: Various README versions data performance.
In Set-2, we focus on the efficacy of different data versions when finetuning with the GPT-2 model. The results, as reflected in Table 2, indicate that high-quality expert data (v2) demonstrates clear superiority over unexamined expert data (v1), emphasizing the crucial role of ECE-examiner in enhancing data quality. Furthermore, high-quality synthetic data (v5) outperforms the unexamined expert data (v1), underscoring the significant value of data augmentation in the ECE pipeline. Notably, the combination of v2 and v5 shows improved performance over v2 alone, suggesting that including synthetic data is beneficial. This composite approach of v2+v5 leading the rank underscores the efficacy of our EAE Human-AI team.
5.3.2 Expert and Synthetic Data Integration
ROUGE1 J2L J+C2L J+G2L J+C+G2L
RANDOM 19.03 19.65 26.21 26.97
SYNTAX(R) 19.82(-2.28) 20.65(-2.78) 27.56(-8.36) 28.33(-8.84)
SEMANTIC(R) 19.74(-1.11) 19.94(-1.39) 26.42(-0.92) 27.19(-1.96)
MODEL(R) 20.18(-0.8) 20.46(-1.88) 27.9(-2.37) 28.65(-4.22)
UMLS-f J2L J+C2L J+G2L J+C+G2L
RANDOM 8.05 9.09 15.78 16.72
SYNTAX(R) 8.53(-0.11) 9.99(-1.00) 17.76(-5.11) 17.42(-5.54)
SEMANTIC(R) 8.98(-0.38) 9.07(-0.69) 15.93(-0.02) 16.74(-0.93)
MODEL(R) 9.18(-0.63) 9.4(-0.78) 18.34(-2.5) 18.45(-2.55)
Table 3: Different data selection methods performance.
In Set-3, we explored the effects of various data selection strategies on data augmentation outcomes. We found that, compared to the other two methods, the results of SEMANTIC are closer to RANDOM. Also, the SEMANTIC-based ranking remained fairly consistent between SEMANTIC and SEMANTIC_r. However, significant differences were observed in SYNTAX and SYNTAX_r, MODEL and MODEL_r. Table 3 highlights two main findings: firstly, data selection is crucial, as all methods have better performance with higher-ranked data (e.g., SYNTAX, SEMANTIC, MODEL) over lower (e.g., SYNTAX_r, SEMANTIC_r, MODEL_r). All three methods, SYNTAX, SEMANTIC, MODEL, have better results than the RANDOM baseline. Secondly, the effectiveness of a selection algorithm is proportional to the range it creates between the top and last ranking items, with a larger gap indicating a more effective selection process. This implies that with an equal volume of v5 data, the selection methods prioritize data that maximizes the benefits of augmentation and relegates the least impactful data to the lower echelon.
5.3.3 Retrieval-augmented Generation
ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1
Without data augmentation (v2)
jargon2lay 19.47 6.06 18.53 14.74 8.76
jargon+context2lay 19.40 6.40 18.38 15.12 9.24
jargon+gen2lay 26.99 10.76 25.57 21.19 16.88
jargon+context+gen2lay 27.58 11.31 26.35 21.72 17.12
With data augmentation (v2+v5 with SYNTAX)
jargon2lay 21.98 7.42 20.88 16.98 10.95
jargon+context2lay 22.13 7.90 21.04 17.54 10.71
jargon+gen2lay 29.82 13.14 28.47 24.42 20.09
jargon+context+gen2lay 29.89 13.48 28.49 24.65 20.27
Table 4: Efficacy of incorporating medical documentation context and general definition in input data.
Set-4 results underscore the significant improvement in model performance when input data is enriched with UMLS-retrieved general definitions. As illustrated in Table 4, regardless of whether we utilize only expert-annotated data or data augmentation with AI-synthetic data, including general definitions consistently enhances effectiveness. This finding confirms the value of RAG with the general definition in the lay definition generation task. Meanwhile, adding context to the input data yields a moderate impact on model performance.
5.3.4 Model Performances Against ChatGPT
In Set-5, we observed that LLAMA2-7B’s ROUGE-1 and UMLS-F1 metrics surpassed GPT-3.5-turbo in the jargon2lay task post-training. For the jargon+gen2lay setting, DistilGPT2-88M demonstrated equivalent results to GPT-3.5-turbo, while BioGPT’s performance exceeded it, and LLAMA2-7B significantly outperformed the GPT-3.5-turbo. These findings, as depicted in Figure 5, emphasize the effectiveness of open-source, mobile-adapted smaller models when appropriately fine-tuned with high-quality datasets, offering a promising avenue for deploying lightweight yet powerful NLP tools in mobile healthcare applications to help patient education.
6 Human Evaluation
6.1 Human Evaluation settings
Our human evaluation was conducted by 5 human evaluators  [^10] All of whom hold bachelor’s degrees. We randomly selected 50 pairs of (jargon, lay definitions) from the test dataset for this human evaluation. The task for evaluators was to reference the expert definitions and choose a binary preference among the following four groups of definitions: 1) DistilGPT2-J2L vs. GPT-3.5-turbo, 2) DistilGPT2-J+C+G2L vs. GPT-3.5-turbo, 3) LLAMA2-J2L vs. GPT-3.5-turbo, 4) LLAMA2-J+C+G2L vs. GPT-3.5-turbo.
6.2 Human Evaluation Results
Figure 6: Human evaluation results (win rate).
As shown in Figure 6, although the result of adding context and general definition (DistilGPT2-J+C+G2L) is better than DistilGPT2-J2L, the win rate of the two DistilGPT2 models’ output is still significantly behind the result of GPT-3.5-turbo. For LLAMA-2, generating lay definitions directly from jargon is still not as good as GPT-3.5-turbo, but adding context and general definitions is of great help. Human evaluators prefer LLAMA2-J+C+G2L more than GPT-3.5-turbo. There are some inconsistencies between the results of human evaluation and automatic evaluation. We further interviewed two medical students  [^11] Both have hospital internship experience and concluded the following conclusions to help future improvements: 1. While all our in-house systems perform satisfactorily, GPT-3.5-turbo stands out for its flexibility and user-friendliness. Particularly, it excels in elaborating complex medical terms, offering detailed explanations and practical examples for better comprehension. 2. Recent advancements Cai et al. (2023); Zhang et al. (2023b) reveal ChatGPT’s role in enhancing patient education through interactive formats like NoteAid-interactive Zhang et al. (2023a). It enables patients to actively ask questions and seek clarifications, while the AI tailors responses to aid their understanding. This interactive approach, absent in traditional dictionary-style definitions like our README dataset, calls for next-step model distillation work or further refinement in aligning the in-house system’s outputs with patient preferences. Additionally, the development of automatic metrics aligning closely with human evaluation is another critical next step.
7 Limitations and Ethical Considerations
8 Conclusions
In conclusion, our study underscores the potential of NLP to democratize medical knowledge, enabling patient-centric care by simplifying complex medical terminology. By crafting the README dataset and integrating a data-centric Human-AI collaboration, we have not only enriched the dataset quality but also broadened the horizons for training AI models in resource-constrained scenarios. Our findings reveal that even smaller, open-source models can rival the capabilities of advanced, proprietary systems like ChatGPT when fine-tuned with meticulously curated data. This research paves the way for innovative patient education tools, making strides toward a future where all patients can navigate their health information with ease and understanding.
@@ 8 Conclusions
Examiner - 1
model = text-davinci-003(GPT - 3)
[This examiner prompt does not use context as discussed in section 3.2.1]
Decide whether the general definition is correct.
If we can generate the lay definition from the general definition then answer is yes.
term : mg
general definition : this is short for milligram which is 1/1000 of a gram usually considered a small amount.
lay definition : A tiny amount of something, usually a drug.
answer : yes
term : vitamin c
general definition : [‘A nutrient that the body needs in small amounts to function and stay healthy. Vitamin C helps fight infections, heal wounds, and keep tissues healthy. It is an antioxidant that helps prevent cell damage caused by free radicals (highly reactive chemicals). Vitamin C is found in all fruits and vegetables, especially citrus fruits, strawberries, cantaloupe, green peppers, tomatoes, broccoli, leafy greens, and potatoes. It is water-soluble (can dissolve in water) and must be taken in every day. Vitamin C is being studied in the prevention and treatment of some types of cancer.’]
lay definition : A nutrient needed by the body to form and maintain bones, blood vessels, and skin.
answer : yes
term : nodule
general definition : [‘A small lump, swelling or collection of tissue.’]
lay definition : A growth or lump that may be cancerous or not.
answer : yes
term : qd
general definition : [‘Occurring or done each day.’]
lay definition : Every day.
answer : yes
If the general definition contains many words from the term then answer is no.
term : prochlorperzine
general definition : [‘prochlorperzine’,  ’]
lay definition : A drug used to prevent or reduce nausea and vomiting.
answer : no
term : mg
general definition : [‘mg’]
lay definition : A tiny amount of something, usually a drug.
answer : no
If the lay definition can not be generated by the general definition then answer is no.
term : Virt - Vite
general definition : [‘Virt’,  - ’, The determination of the amount of Vitamin E present in a sample.’]
lay definition : A mix of vitamins. It provides vitamin B-6, vitamin B-12 and folic acid to people who do not have enough of these for good health.
answer : no
Augmenter
system_prompt = "your job is to generate a general definition of the term for a professional medical student"
model="gpt-3.5-turbo(ChatGPT API)",
messages=[
        {"role": "system", "content": system_prompt},
        {"role": "user", "content": ""},
        {"role": "user", "content": "term : incisional."},
        {"role": "assistant", "content": "general definition : An intentional cut made to an individuals body with the intent of performing a diagnostic or therapeutic intervention."},
        {"role": "user", "content": "term : PO"},
        {"role": "assistant", "content": "general definition : Of, or relating to, or affecting, or for use in the mouth.."},
        {"role": "user", "content": prompt_t}
Examiner - 2
model = text-davinci-003(GPT - 3)
exactly same as that of Examiner - 1

### v2

@@ Appendix D Data Quality Checking and Train/Eval/Test Split after EAE Pipeline
After the end of the EAE, we use R-exp_good and R-syn_good as high-quality data for our system. The dataset was split into two categories: human examination data (which will also be used as final evaluation and test data since medical experts examine this split), and training data, where we ensure the medical jargon in the human examination will not appear in the training split. We sampled 500 medical jargon terms each from the R-exp_good and R-syn_good datasets. Therefore the human examination split consisted of 1000 unique terms, each accompanied by general and lay definitions to be rated based on two criteria:
@@
Although the lay definition is not incorrect, it is very wordy and complex and matches less to the lay definition. We consider this soft correlated and not hard correlated. A Hard Correlation automatically implies a Soft Correlation.
Two medical students  [^15] Both have hospital internship experience help us finish this human examination. Our findings revealed that 88% of R-exp_good and 100% of R-syn_good met the Hard Correlation criteria. There is a 100% soft Correlation criteria for both R-exp_good and R-syn_good. After correcting individual invalid data (e.g., those cases where Soft is not satisfied), we used this human examination dataset as evaluation and test data (in a 1:1 ratio).
@@
Appendix E Factuality metrics: UMLS-F1
The assessment of factual accuracy in generated lay definition leverages the UMLS concept overlap metric. The Unified Medical Language System (UMLS), established by Bodenreider (2004), significantly contributes to the biomedical domain’s interoperability. It achieves this by amalgamating and disseminating a comprehensive collection of biomedical terminologies, classification systems, and coding standards from many sources. By doing so, UMLS aids in reconciling semantic variances and representational disparities found across different biomedical concept repositories.
For the identification and alignment of medical named entities within texts to their corresponding biomedical concepts in UMLS, we employed the Scispacy library  [^16] We used the Scispacy en_core_sci_lg model.. Scispacy excels in identifying and clarifying entities, thus facilitating the accurate association of named entities found in lay definitions with the relevant UMLS concepts. This capability is critical for evaluating the lay definitions’ factual accuracy and is used by recent related work Adams et al. (2023).
The analytical process for lay definitions utilizes metrics of precision and recall. Precision represents the ratio of concepts present in both the generated and reference lay definitions, serving as a measure of the generated lay definition’s factual correctness. In contrast, recall evaluates how well the information in the generated lay definition matches the intended content, reflecting the relevance of the presented information.
To calculate these metrics, we consider the concept sets from both the reference lay definition ($C_{ref}$) and the generated lay definition ($C_{gen}$). The formulas for recall and precision are as follows:
$\text{Recall}=\frac{|C_{ref}\cap C_{gen}|}{|C_{ref}|}$
$\text{Precision}=\frac{|C_{ref}\cap C_{gen}|}{|C_{gen}|}.$
The F1 score, derived from the above precision and recall values, is reported to provide a balanced measure of the generated lay definition’s accuracy and relevance.
Appendix F More Experimental Settings
We use the following base models: GPT-2  [^17] https://huggingface.co/gpt2 Radford et al. (2019), DistilGPT2  [^18] https://huggingface.co/distilgpt2, BioGPT  [^19] https://huggingface.co/microsoft/biogpt Luo et al. (2022a), and Llama2  [^20] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023) in our experiments. We trained the base models on the different README dataset variants with Supervised Fine-tuning for 100000 steps (batch size 8)  [^21] We did all the experiments with 1 NVIDIA Tesla RTX 8000 GPU - 40 GB memory, with Adam optimizer – betas=(0.9,0.999), epsilon=1e-08, learning rate=5e-04.. In all our evaluations, we used a beam size of 4, no-repeat-ngram-size=2, and minimum length and maximum length of sentences were set as (10, 100). We used five different random seeds to sample training data for all our experiments, and the scores reported in the tables are the average of these random seeds.
Here, we provide an example for jargon2lay(J2L), jargon+context2lay(J+C2L), jargon+gen2lay(J+G2L), and jargon+context+gen2lay(J+C+G2L) for easier understanding. Let’s assume we one data point in README dataset with (jargon, EHR context, lay def, general def) format:
  • jargon: EGD
  • lay def: [esophagogastroduodenoscopy] A procedure that looks at the food pipe, stomach, and the first part of the small bowel.
  • EHR context: [ * * 11 - 22 * * ] EGD Grade I varices - ablated [ * * 11 - 22 * * ] sigmoidoscopy friability , reythema , congest and abnormal vasularity in a small 5 mm area of distal rectum .
  • general def: [’An endoscopic procedure that visualizes the upper part of the gastrointestinal tract up to the duodenum.’]
The input of different settings J2L, J+C2L, J+G2L, J+C+G2L can be found on Table 7.
jargon2lay(J2L):
jargon term: EGD
lay definition:
jargon+context2lay(J+C2L):
In this task, we ask for your expertise in generating the corresponding lay definition from the medical jargon. Mainly, we provide the target medical jargon term along with the contextual snippets in which they appear in the text. We need you to generate a lay definition for this jargon term.
jargon term: EGD
context: [ * * 11 - 22 * * ] EGD Grade I varices - ablated [ * * 11 - 22 * * ] sigmoidoscopy friability , reythema , congest and abnormal vasularity in a small 5 mm area of distal rectum .
lay definition:
jargon+gen2lay(J+G2L):
In this task, we ask for your expertise in generating the corresponding lay definition from the medical jargon. Mainly, we provide the target medical jargon term. In addition, we also provide a definition from the dictionary for reference. We need you to generate a lay definition for this jargon term.
jargon term: EGD
dictionary definition: [’An endoscopic procedure that visualizes the upper part of the gastrointestinal tract up to the duodenum.’]
lay definition:
jargon+context+gen2lay(J+C+G2L):
In this task, we ask for your expertise in generating the corresponding lay definition from the medical jargon. Mainly, we provide the target medical jargon term along with the contextual snippets in which they appear in the text. In addition, we also provide a definition from the dictionary for reference. We need you to generate a lay definition for this jargon term.
jargon term: EGD
context: [ * * 11 - 22 * * ] EGD Grade I varices - ablated [ * * 11 - 22 * * ] sigmoidoscopy friability , reythema , congest and abnormal vasularity in a small 5 mm area of distal rectum .
dictionary definition: [’An endoscopic procedure that visualizes the upper part of the gastrointestinal tract up to the duodenum.’]
lay definition:
Table 7: The prompt of different settings J2L, J+C2L, J+G2L, J+C+G2L.
README-EAE
       inputs : README-exp
       output : Final README dataset with best general definitions and lay definitions
       $Initialize:$ $exam_{p}=examinerpromptforgpt3$;
       $aug_{p}=augmenterpromptsforChatGPT$;
       foreach $datapoint:README-exp$ do
            if $exam_{p}(datapoint)==yes$ then
                   R-exp_good.add(datapoint);
             end if
            else
                   R-exp_bad.add(datapoint);
             end if
       end foreach
      foreach $datapoint:R-exp\_bad$ do
            $temp=aug_{p}(datapoint)$;
             R-syn.add(temp);
             if $exam_{p}(temp)==yes$ then
                   R-syn_good.add(temp);
             end if
            else
                   R-syn_bad.add(temp);
             end if
       end foreach
      return R-syn_bad;
Algorithm 1 Algorithm for Data Cleaning
@@ Appendix F More Experimental Settings
Examiner - 1 model = "gpt-3.5-turbo(ChatGPT API)" [This examiner prompt does not use context as discussed in section 3.2.1] Decide whether the general definition is correct. If we can generate the lay definition from the general definition then answer is yes. term : mg general definition : this is short for milligram which is 1/1000 of a gram usually considered a small amount. lay definition : A tiny amount of something, usually a drug. answer : yes term : vitamin c general definition : [‘A nutrient that the body needs in small amounts to function and stay healthy. Vitamin C helps fight infections, heal wounds, and keep tissues healthy. It is an antioxidant that helps prevent cell damage caused by free radicals (highly reactive chemicals). Vitamin C is found in all fruits and vegetables, especially citrus fruits, strawberries, cantaloupe, green peppers, tomatoes, broccoli, leafy greens, and potatoes. It is water-soluble (can dissolve in water) and must be taken in every day. Vitamin C is being studied in the prevention and treatment of some types of cancer.’] lay definition : A nutrient needed by the body to form and maintain bones, blood vessels, and skin. answer : yes term : nodule general definition : [‘A small lump, swelling or collection of tissue.’] lay definition : A growth or lump that may be cancerous or not. answer : yes term : qd general definition : [‘Occurring or done each day.’] lay definition : Every day. answer : yes If the general definition contains many words from the term then answer is no. term : prochlorperzine general definition : [‘prochlorperzine’, ’] lay definition : A drug used to prevent or reduce nausea and vomiting. answer : no term : mg general definition : [‘mg’] lay definition : A tiny amount of something, usually a drug. answer : no If the lay definition can not be generated by the general definition then answer is no. term : Virt - Vite general definition : [‘Virt’, - ’, The determination of the amount of Vitamin E present in a sample.’] lay definition : A mix of vitamins. It provides vitamin B-6, vitamin B-12 and folic acid to people who do not have enough of these for good health. answer : no
Augmenter system_prompt = "your job is to generate a general definition of the term." model="gpt-3.5-turbo(ChatGPT API)", messages=[ {"role": "system", "content": system_prompt}, {"role": "user", "content": ""}, {"role": "user", "content": "term : incisional."}, {"role": "assistant", "content": "general definition : An intentional cut made to an individuals body with the intent of performing a diagnostic or therapeutic intervention."}, {"role": "user", "content": "term : PO"}, {"role": "assistant", "content": "general definition : Of, or relating to, or affecting, or for use in the mouth.."}, {"role": "user", "content": prompt_t} ]
Examiner - 2 model = "gpt-3.5-turbo(ChatGPT API)" exactly same as that of Examiner - 1